A statistician determines that a certain baseball player (Player A) managed to hit a home run in 30% of the games he played. It is also found that when player A hits a home run in a game, his team's chances of winning are 80%. What is the probability that the player hits a home run, and his team does not win the game?
Answer by ikleyn(52937)

.
A statistician determines that a certain baseball player (Player A) managed to hit
a home run in 30% of the games he played. It is also found that when player A hits
a home run in a game, his team's chances of winning are 80%. What is the probability
that the player hits a home run, and his team does not win the game?
~~~~~~~~~~~~~~~

It is 0.3*(1-0.8) = 0.3*0.2 = 0.06.    ANSWER

Solved.